A tuning fork vibrating at a frequency of 512 Hz is held over the top of a jar filled with water and fitted with a tap at the bottom. If the jar is 60 cm high and the speed of sound is 350 m/s, describe what occurs as the jar slowly emptied.
1
Expert's answer
2020-05-25T11:19:15-0400
"\lambda=\frac{v}{f}=\frac{350}{512}=0.684\ m"

First resonance:


"l_1=\frac{\lambda}{4}=\frac{0.684}{4}=0.17\ m=17\ cm"

"l_2=3\frac{\lambda}{4}=3\frac{0.684}{4}=0.51\ m=51\ cm"

As the jar slowly emptied we can hear resonance when the height of the jar would be 17 cm and 51 cm.